# A subset $S$ of vertices in a planar graph $G$ is a free set if, for every set $P$ of $|S|$ points in the plane, there exists a straight-line crossing-free drawing of $G$ in which vertices of $S$ are mapped to distinct points in $P$. In this survey, we review

### v1

@@
Free Sets in Planar Graphs: History and Applicationsthanks: This research is partially funded by the grants from NSERC and the University Research Chair of University of Ottawa.
@@
  • results on the existence of large free sets in planar graphs and subclasses of planar graphs,
@@ 1 Introduction
In 2005, an online game called Planarity became wildly popular. In this single-player game, created by John Tantalo based on a concept by Mary Radcliffe, the player is presented with a straight-line drawing $\Gamma$ of a planar graph $G$. Although the graph $G$ is planar, the drawing $\Gamma$ is not crossing-free; some pairs of edges cross each other and an edge may even contain a vertex in its interior. Since $\Gamma$ draws the edges of $G$ as straight line segments, the drawing is completely determined by the vertex locations. The job of the player is to move vertices in the drawing $\Gamma$ in order to obtain a crossing-free drawing of $G$. Since $G$ is planar, Fáry’s Theorem guarantees that this always possible.
@@
The act of moving some vertices in $\Gamma$ to obtain a crossing-free drawing is referred to as untangling. For a combinatorially inclined player, this naturally leads to the question: What is the fewest number, $\operatorname{shift}(\Gamma)$, of vertices that need to be moved in order untangle $\Gamma$? [^1] The Planarity game actually measures the amount of time the player takes to accomplish this task, not the number of vertices moved, but we won’t let facts get in the way of a good story. Equivalently, what is the maximum number, $\operatorname{fix}(\Gamma):=|V(G)|-\operatorname{shift}(\Gamma)$ of vertices that can be kept fixed while untangling $G$? For a planar graph $G$, let $\operatorname{fix}(G):=\min\{\operatorname{fix}(\Gamma):\text{$\Gamma$ is a straight-line drawing of $G$}\}$. For a family $\mathcal{F}$ of planar graphs, this defines a function $\operatorname{fix}_{\mathcal{F}}(n):=\min\{\operatorname{fix}(G):G\in\mathcal{F},\,|V(G)|=n\}$. [^2] Technically $\operatorname{fix}_{\mathcal{F}}(n)$ is undefined if there are no $n$-vertex graphs in the family $\mathcal{F}$. We will ignore this, since all the graph families we consider are infinite and have at least one $n$-vertex member for each $n\in\mathbb{N}$.
@@
Determining the asymptotic growth of $\operatorname{fix}_{\mathcal{F}}(n)$ turns out to be a challenging problem, even for very simple classes $\mathcal{F}$. In 1998, Mamoru Watanabe had already asked this problem for the class $\mathcal{C}$ of cycles and we now know that $\operatorname{fix}_{\mathcal{C}}(n)\in\Omega(n^{2/3})\cap O((n\log n)^{2/3})$ but neither the lower bound, due to Pach and Tardos [46], nor the upper bound, due to Cibulka [23], is easy to prove. For the class $\mathcal{G}$ of all planar graphs, it is only known that $\operatorname{fix}_{\mathcal{G}}(n)\in\Omega(n^{1/4})\cap O(n^{0.4948})$ [17, 42, 38, 20], despite the problem being studied since Pach and Tardos [46] asked for a polynomial lower bound over $20$ years ago.
@@
In this survey, we study various kinds of vertex subsets of a planar graph $G$: proper-good sets, collinear sets, free-collinear sets, and free sets. Each of these definitions is, at first glance, more stringent than the one that precedes it. Indeed, it follows immediately from definitions that every free set is a free-collinear set, every free-collinear set is a collinear set, and every collinear set is a proper-good set. The definition of a proper-good set is the most relaxed, which makes it easy to find large proper-good sets, even by hand using a pencil and a paper that contains a crossing-free drawing of $G$. At the other extreme, free sets satisfy a very strong property that says we can place the vertices of a free set on any given pointset and then find locations for the other vertices of $G$ that results in a non-crossing straight-line drawing of $G$. This makes free sets incredibly useful for many applications, including untangling.
@@
A key result in this area, discovered over a period of roughly 15 years, is that all of these definitions are equivalent: A vertex subset in a planar graph is proper-good if and only if it is collinear if and only if it is free-collinear if and only if it is free. Meanwhile, proper-good sets, collinear sets, free-collinear sets, and free sets were being used to resolve problems in graph drawing and related areas. With the benefit of hindsight, we attempt to organize and present two decades of results on free sets and their applications in a way that makes them as easy to understand and as useful as possible.
@@
The rest of this survey is structured as follows: In Section 2 we present four definitions of free sets and explain why these four definitions are equivalent. In Section 3 we present upper and lower bounds on the size of free sets in planar graphs and various subclasses of planar graphs. In Section 4 we describe applications of free sets to a variety of graph drawing problems. Section 6 concludes with a list of open problems and directions for further research.
@@ 2 Four Definitions of Free Sets
For definitions of standard graph theoretic terms and notations used in this survey (such a treewidth, $k$-trees, independent set, etc.) the reader is referred to the textbook by Diestel [26].
@@
A drawing of a graph is representation of the graph in which each vertex $v$ is represented by a distinct point in the plane and each edge $vw$ is represented by a simple open curve in the plane the closure of which has $v$ and $w$ as endpoints. When discussing a particular drawing, we do not distinguish between a vertex and the point that represents it, or an edge and the curve that represents it. A drawing is a straight-line drawing if each edge is an open line segment. In a drawing, two edges may have a non-empty intersection, and an edge may contain a vertex that is not one of its endpoints. A drawing is a crossing-free drawing if this does not occur: any two distinct edges are disjoint and no vertex is contained in any edge. [^3] What we call a crossing-free drawing is sometimes called a plane drawing or an embedding (in the plane). What we call a straight-line crossing-free drawing is sometimes called a geometric embedding or Fáry embedding.
@@
Recall that for any $n$-vertex planar graph $G_{0}$, on at least $3$ vertices, it is possible to add edges to $G_{0}$ to obtain a planar graph $G$ with $3(n-2)$ edges. The resulting graph $G$ is a triangulation: in any crossing-free drawing of this graph, each of the faces has exactly three edges on its boundary. A near-triangulation is a biconnected embedded graph whose inner faces are all triangles, i.e., bounded by three edges. The dual $G^{*}$ of an embedded graph $G$ is a graph whose vertices are the faces of $G$ that contains an edge between two vertices $f$ and $g$ if the corresponding faces of $G$ have a common edge on their boundary. For a planar graph $G$ with outer face $f_{0}$, the weak dual $G^{+}$ of $G$ is defined as $G+:=G^{*}-f_{0}$, the graph obtained from the dual $G^{+}$ by removing the vertex corresponds to the outer face $f_{0}$ of $G$. A graph is outerplanar if it has a crossing-free drawing with all its vertices on the outer face. A graph equipped with such a crossing-free drawing is an outerplane graph. A chord of a cycle $C$ is an edge that is not part of $C$ but both of its endpoints are vertices of $C$.
@@
Under our definitions, a planar graph is a graph that has a crossing-free drawing. Fáry’s Theorem [58, 35, 53] asserts that a graph is planar if and only if it has a straight-line crossing-free drawing. Today, we use Fáry’s Theorem without thinking about it, but it is worthwhile spending a moment reflecting on what would happen if Fáry’s Theorem were not true. There would be two kinds of planar graphs, topological and straight-line, with the latter being a special case of the former. Any result proven for straight-line planar graphs would need a separate proof for topological planar graphs. Edge contractions, commonly used in inductive proofs for planar graphs would be off-limits for proofs about straight-line planar graphs. In order to establish the existence of a straight-line planar graph with some property one would need to find a topological planar graph with that property and then show that it has a straight-line crossing-free drawing.
@@
Let $\Gamma$ be a crossing-free drawing of a planar graph $G$. A simple closed curve [^4] A simple closed curve $C$ is a continuous function $C:[0,1]\to\mathbb{R}^{2}$ such that $C(0)=C(1)$ and $C(a)\neq C(b)$ for all $10\leq a<b<$ 1. $C:[0,1]\to\mathbb{R}^{2}$ is good with respect to $\Gamma$ if $C(0)$ is not contained in any edge or vertex of $\Gamma$. $C$ is proper with respect to $\Gamma$ if, for each edge $e$ of $\Gamma$, the intersection of $C$ and the closure of $e$ is either empty, consists of a single point (possibly an endpoint of $e$), or consists of the entire edge $e$. In essence, a curve that is proper-good with respect to $\Gamma$ behaves the way a line would behave if $\Gamma$ were a straight-line crossing-free drawing.
@@
Let $S:=(v_{1},\ldots,v_{s})$ be an ordered subset of vertices in a planar graph $G$.
@@
    $S$ is a proper-good set [24] if there exists a crossing-free drawing $\Gamma$ of $G$, a curve $C:[0,1)\to\mathbb{R}^{2}$ that is proper-good with respect to $\Gamma$, and $10<x_{1}<\cdots<x_{s}<$ 1 such that $C(x_{i})$ is the location of $v_{i}$ in $\Gamma$, for each $i\in\{1,\ldots,s\}$. In other words, we encounter the vertices of $S$, in order, while traversing $C$.
@@
We note that our definitions deviate from much of the literature by treating $S$ as an ordered set. Later, we may say that some (unordered) vertex subset $S$ of $G$ is a proper-good, collinear, free-collinear, or free set. In these cases, we mean that there is some permutation of $S$ that defines an ordered set that satisfies the relevant definition. When we want to emphasize the difference between these two we will say that (an ordered set) $S$ is an ordered free set or that (a set) $S$ is an unordered free set. This distinction becomes important when discussing two or more graphs with the same vertex set. A set $S$ of vertices may be an unordered free set in two graphs $G_{1}$ and $G_{2}$ but there may be no permutation of $S$ that is an ordered free set in $G_{1}$ and in $G_{2}$.
A few properties of free sets are immediate from the definition and we will use them throughout. If $S$ is an ordered free set in a graph $G$ then:
@@
Unordered collinear sets and unordered free-collinear sets were defined first by Ravsky and Verbitsky [48] and were implicit in [17]. In fact, Ravsky and Verbitsky [48] posed the equivalence between (unordered) collinear sets and (unordered) free-collinear set as an open question, and conjectured a negative answer. Thus 1 below disproves that conjecture. Proper-good sets (as unordered sets) were introduced first by Da Lozzo et al. [24].
@@
The proof of Lemma 1, below, gives one example of why definitions based on ordered sets are useful. Another example, discussed in Section 4, is the problem of simultaneous embedding with mapping.
@@
Suppose $(S,\prec)$ contains a chain $v_{1}\prec\cdots\prec v_{\ell}$ of length $\ell\geq\sqrt{k}$. Then $(v_{1},\ldots,v_{\ell})$ is a free set since it is a subsequence of $S$. For each $i\in\{1,\ldots,\ell\}$, let $(x_{i},y_{i})$ be the coordinates given to $v_{i}$ in $\Gamma$. By the definition of free set, $G$ has a straight-line crossing-free drawing in which $v_{i}$ is mapped to $(x_{i},y_{i})$, for each $i\in\{1,\ldots,\ell\}$. Therefore, $\operatorname{fix}(\Gamma)\geq\ell\geq\sqrt{k}$ for any straight-line drawing $\Gamma$ of $G$. Therefore $\operatorname{fix}(G)\geq\sqrt{k}$.
[…]
For most classes $\mathcal{X}$ of planar graphs, including trees, outerplanar graphs, and all planar graphs, the best known bounds (and for some classes asymptotically optimal bounds) for $\operatorname{fix}_{\mathcal{X}}(n)$ can be obtained by an application of Lemma 1, along with a result on free sets in graphs in the class $\mathcal{X}$ [38, 24, 17, 48]. The sole exception is the class $\mathcal{C}$ of cycles, for which Cibulka [23] uses repeated applications of the Erdős-Szekeres Theorem to obtain a lower bound of $\Omega(n^{2/3})$.
@@
    $S$ is a (ordered) proper-good set.
@@
    $S$ is a (ordered) free-collinear set.
@@
    $S$ is a (ordered) free set.
Note that, unlike the other three definitions, the definition of proper-good set is purely combinatorial/topological; it has no mention of straight lines or vertex coordinates. In this respect, 1 plays a role similar to Fáry’s Theorem. Using the proper-good set definition allows purely combinatorial methods to be used in the search for free sets. For any proper-good set there is a short easily-verifiable certificate (a crossing-free drawing of $G$ and a description of the curve $C$) that $S$ is indeed a proper-good set. In contrast, in order to show directly that $S$ is a free-collinear set (or a free set) involves proving a fact that must hold for all choices of $x_{1},\ldots,x_{k}$ (or $(x_{1},y_{1}),\ldots,(x_{k},y_{k})$, respectively). Working with this definition directly would involve (at the very least) a significant amount of linear algebra.
The proof of 1 is spread across a few different works [24, 29, 17], which we describe below. First we remark that it is immediate from the definitions that every free set is a free-collinear set and every free-collinear set is a free set. It is also easy to see that every collinear set is a proper-good set by taking a straight-line crossing-free drawing $\Gamma$ of $G$ with the vertices of the (ordered) collinear set $S$ on the ${x}$-axis in increasing order of ${x}$-coordinate. Let $s$ be a line segment that is contained in the ${x}$-axis and that contains every intersection between $\Gamma$ and the ${x}$-axis in its interior. Then $s$ is proper and good with respect to $\Gamma$. Finally, make $s$ into a closed curve by joining the endpoints of $s$ with a curve $C$ that completely avoids $\Gamma$.
@@
The proof that every proper-good set is a collinear set is due to Da Lozzo et al. [24] and is illustrated in Fig. 1. For each edge of $G$ that crosses $C$, the authors introduce a dummy vertex (white vertices in Fig. 1(a)) that splits the edge into two edges with a common vertex on $C$. The resulting graph $G^{\prime}$ then has three types of vertices: a set $X$ of inner vertices (contained in the interior of $C$), a set $Y$ of boundary vertices (contained in $C$) and a set $Z$ of outer vertices (the vertices in the exterior of $C$). These steps are depicted in Fig. 1(a).
Refer to caption Refer to caption

### v2

@@
Free Sets in Planar Graphs: History and Applicationsthanks: This research is partially funded by the grants from NSERC and University of Ottawa Research Chair.
@@ 1 Introduction
In 2005, an online game called Planarity became wildly popular [^1] https://en.wikipedia.org/wiki/Planarity. In this single-player game, created by John Tantalo based on a concept by Mary Radcliffe, the player is presented with a straight-line drawing $\Gamma$ of a planar graph $G$. Although the graph $G$ is planar, the drawing $\Gamma$ is not crossing-free; some pairs of edges cross each other, and an edge may even contain a vertex in its interior. Since $\Gamma$ draws the edges of $G$ as straight-line segments, the drawing is completely determined by the vertex locations. The job of the player is to move vertices in the drawing $\Gamma$ in order to obtain a crossing-free drawing of $G$. Since $G$ is planar, Fáry’s Theorem guarantees that this is always possible.
@@
The act of moving some vertices in $\Gamma$ to obtain a crossing-free drawing is referred to as untangling. For a combinatorially inclined player, this naturally leads to the question: What is the fewest number, $\operatorname{shift}(\Gamma)$, of vertices that need to be moved in order to untangle $\Gamma$? [^2] The Planarity game actually measures the amount of time the player takes to accomplish this task, not the number of vertices moved. Equivalently, what is the maximum number, $\operatorname{fix}(\Gamma):=|V(G)|-\operatorname{shift}(\Gamma)$ of vertices that can be kept fixed while untangling $G$? For a planar graph $G$, let $\operatorname{fix}(G):=\min\{\operatorname{fix}(\Gamma):\text{$\Gamma$ is a straight-line drawing of $G$}\}$. For a family $\mathcal{F}$ of planar graphs, this defines a function $\operatorname{fix}_{\mathcal{F}}(n):=\min\{\operatorname{fix}(G):G\in\mathcal{F},\,|V(G)|=n\}$. [^3] Technically $\operatorname{fix}_{\mathcal{F}}(n)$ is undefined if there are no $n$-vertex graphs in the family $\mathcal{F}$. We will ignore this, since all the graph families we consider are infinite and have at least one $n$-vertex member for each $n\in\mathbb{N}$.
@@
Determining the asymptotic growth of $\operatorname{fix}_{\mathcal{F}}(n)$ turns out to be a challenging problem, even for very simple classes $\mathcal{F}$. In 1998, Mamoru Watanabe had already asked this problem for the class $\mathcal{C}$ of cycles and we now know that $\operatorname{fix}_{\mathcal{C}}(n)\in\Omega(n^{2/3})\cap O((n\log n)^{2/3})$ but neither the lower bound, due to Pach and Tardos [48], nor the upper bound, due to Cibulka [23], is easy to prove. For the class $\mathcal{G}$ of planar graphs, the problem has been studied since Pach and Tardos [48] asked for a polynomial lower bound over $20$ years ago. The best known bounds to date are $\operatorname{fix}_{\mathcal{G}}(n)\in\Omega(n^{1/4})\cap O(n^{0.4948})$ [17, 44, 39, 20].
@@
In this survey, we study various kinds of vertex subsets of a planar graph $G$: proper-good sets, collinear sets, free-collinear sets, and free sets. Each of these definitions is, at first glance, more stringent than the one that precedes it. Indeed, it follows immediately from definitions that every free set is a free-collinear set, every free-collinear set is a collinear set, and every collinear set is a proper-good set. The definition of a proper-good set is the most relaxed, which makes it easy to find large proper-good sets, even by hand using a pencil and a paper that contains a crossing-free drawing of $G$. At the other extreme, free sets satisfy a very strong property that says we can place the vertices of a free set on any given pointset and then find locations for the other vertices of $G$ that result in a non-crossing straight-line drawing of $G$. This makes free sets incredibly useful for many applications, including untangling.
@@
A key result in this area, discovered over roughly 15 years, is that all of these definitions are equivalent: A vertex subset in a planar graph is proper-good if and only if it is collinear, if and only if it is free-collinear, if and only if it is free. Meanwhile, proper-good sets, collinear sets, free-collinear sets, and free sets were being used to resolve problems in graph drawing and related areas. With the benefit of hindsight, we attempt to organize and present two decades of results on free sets and their applications in a way that makes them as easy to understand and as useful as possible.
@@
The rest of this survey is structured as follows: In Section 2 we present four definitions of free sets and explain why these four definitions are equivalent. In Section 3 we present upper and lower bounds on the size of free sets in planar graphs and various subclasses of planar graphs. In Section 4 we describe applications of free sets to a variety of graph drawing problems. In Section 5, we introduce a one-bend variant of free sets. Section 6 concludes with a list of open problems and directions for further research.
@@ 2 Four Definitions of Free Sets
For definitions of standard graph theoretic terms and notations used in this survey (such a treewidth, $k$-trees, independent set, etc.) the reader is referred to the textbook by Diestel [25]. Throughout this survey, a graph is undirected, has no self-loops, and has no parallel edges, unless specified otherwise.
@@
A drawing of a graph is a representation of the graph in which each vertex $v$ is represented by a distinct point in the plane and each edge $vw$ is represented by a simple open curve in the plane, the closure of which has $v$ and $w$ as endpoints. When discussing a particular drawing, we do not distinguish between a vertex and the point that represents it, or an edge and the curve that represents it. A drawing is a straight-line drawing if each edge is an open line segment. In a drawing, two edges may have a non-empty intersection, and an edge may contain a vertex that is not one of its endpoints. A drawing is a crossing-free drawing if this does not occur: any two distinct edges are disjoint and no vertex is contained in any edge. [^4] What we call a crossing-free drawing is sometimes called a plane drawing or an embedding (in the plane). What we call a straight-line crossing-free drawing is sometimes called a geometric embedding or Fáry embedding.
@@
Recall that for any $n$-vertex planar graph $G_{0}$, on at least $3$ vertices, it is possible to add edges to $G_{0}$ to obtain a planar graph $G$ with $3(n-2)$ edges. The resulting graph $G$ is a triangulation: in any crossing-free drawing of this graph, each of the faces has exactly three edges on its boundary. A near-triangulation is a biconnected embedded graph whose inner faces are all triangles, i.e., bounded by three edges. The dual $G^{*}$ of an embedded graph $G$ is a graph whose vertices are the faces of $G$ that contains an edge between two vertices $f$ and $g$ if these two faces of $G$ have a common edge on their boundaries. For a planar graph $G$ with outer face $f_{0}$, the weak dual $G^{+}$ of $G$ is defined as $G^{+}:=G^{*}-f_{0}$, the graph obtained from the dual $G^{*}$ by removing the vertex corresponding to the outer face $f_{0}$ of $G$. A graph is outerplanar if it has a crossing-free drawing with all its vertices on the outer face. A graph equipped with such a crossing-free drawing is an outerplane graph. A chord of a cycle $C$ is an edge between two vertices of $C$ that are not adjacent in $C$.
@@
Under our definitions, a planar graph is a graph that has a crossing-free drawing. Fáry’s Theorem [59, 35, 54] asserts that a graph is planar if and only if it has a straight-line crossing-free drawing. Today, we use Fáry’s Theorem without thinking about it, but it is worthwhile spending a moment reflecting on what would happen if Fáry’s Theorem were not true. There would be two kinds of planar graphs, topological and straight-line, with the latter being a special case of the former. Any result proven for straight-line planar graphs would need a separate proof for topological planar graphs (assuming it was also true for topological planar graphs). Edge contractions, which obviously preserve topological planarity and are commonly used in inductive proofs for planar graphs, would be off-limits for proofs about straight-line planar graphs.
Indeed, there are a number of questions in which the difference between (not necessarily crossing-free) drawings and straight-line drawings of graphs is profound. A famous example of this is Conway’s Thrackle Conjecture [60], which asserts that if $G$ has a drawing in which each pair of edges share exactly one point (possibly a common endpoint) then the number of edges of $G$ is no more than the number of vertices of $G$. This conjecture has been open since the 1960s but the straight-line version of the question (the Linear Thrackle Problem) has an easy solution, first observed by Erdős [29].
@@
Let $\Gamma$ be a crossing-free drawing of a planar graph $G$. A simple closed curve [^5] A simple closed curve $C$ is a continuous function $C:[0,1]\to\mathbb{R}^{2}$ such that $C(0)=C(1)$ and $C(a)\neq C(b)$ for all $10\leq a<b<$ 1. $C:[0,1]\to\mathbb{R}^{2}$ is good with respect to $\Gamma$ if $C(0)$ is not contained in any edge or vertex of $\Gamma$. The curve $C$ is proper with respect to $\Gamma$ if, for each edge $e$ of $\Gamma$, the intersection of $C$ and the closure of $e$ is either empty, consists of a single point (possibly an endpoint of $e$), or consists of the entire edge $e$. In essence, a curve that is proper-good with respect to $\Gamma$ behaves the way a line would behave if $\Gamma$ were a straight-line drawing.
@@
Let $S:=(v_{1},\ldots,v_{s})$ be an ordered subset (that is, a sequence of distinct elements) of vertices in a planar graph $G$.
@@
    $S$ is a proper-good set [24] if there exists a crossing-free drawing $\Gamma$ of $G$, a curve $C:[0,1]\to\mathbb{R}^{2}$ that is proper and good with respect to $\Gamma$, and $10<x_{1}<\cdots<x_{s}<$ 1 such that $C(x_{i})$ is the location of $v_{i}$ in $\Gamma$, for each $i\in\{1,\ldots,s\}$. In other words, we encounter the vertices of $S$, in order, while traversing $C$.
@@
We note that our definitions deviate from much of the literature by treating $S$ as an ordered set. Later, we may say that some (unordered) vertex subset $S$ of $G$ is a proper-good, collinear, free-collinear, or free set. In these cases, we mean that there is some permutation of $S$ that defines an ordered set that satisfies the relevant definition. When we want to emphasize the difference between these two, we will say that (an ordered set) $S$ is an ordered free set or that (a set) $S$ is an unordered free set. This distinction becomes important when discussing two or more graphs with the same vertex set. A set $S$ of vertices may be an unordered free set in two graphs $G_{1}$ and $G_{2}$, but there may be no permutation of $S$ that is an ordered free set in $G_{1}$ and in $G_{2}$.
A few properties of free sets are immediate from the definition, and we will use them throughout. If $S$ is an ordered free set in a graph $G$, then:
@@
Unordered collinear sets and unordered free-collinear sets were defined first by Ravsky and Verbitsky [50] and also appear implicitly in the work of Bose et al. [17] on untangling. [^6] Although they don’t use this terminology, Bose et al. [17] establish that $\operatorname{fix}_{\mathcal{G}}(n)\in\Omega(n^{1/4})$ by showing that every planar graph has a free collinear set of size $\Omega(n^{1/2})$. In fact, Ravsky and Verbitsky [50] posed the equivalence between (unordered) collinear sets and (unordered) free-collinear sets as an open question, and conjectured a negative answer. 1 below disproves that conjecture. Proper-good sets (as unordered sets) were introduced first by Da Lozzo et al. [24].
@@
Next we turn to the relationship between free sets and the untangling problem in the introduction. In the untangling problem, we are given a straight line drawing $\Gamma$ of $G$ (with crossings) and the goal is to find a large set $F\subseteq V(G)$ and a straight-line crossing-free drawing $\Gamma^{\prime}$ of $G$ such that $\Gamma^{\prime}(v)=\Gamma(v)$ for each $v\in F$. Note that it is not sufficient to choose $F$ to be a free set in $G$ for the following reason: The definition of a free set would only guarantee that we can find a straight-line crossing-free drawing $\Gamma^{\prime}$ such that $\{\Gamma^{\prime}(v):v\in F\}=\{\Gamma(v):v\in F\}$. That is, $\Gamma^{\prime}$ draws the vertices of $F$ on the same point set $P:=\{\Gamma(v):v\in F\}$ in both drawings, but may change the permutation that maps the vertices in $F$ to points in $P$. The following lemma shows that free sets in $G$ are nevertheless useful in obtaining lower bounds on $\operatorname{fix}(G)$. The proof of this lemma also illustrates the usefulness of definitions based on ordered sets. Another example, discussed in Section 4, is the problem of simultaneous embedding with mapping.
@@
By Dilworth’s Theorem, $(S,\prec)$ contains a chain of length at least $\sqrt{k}$ or $(S,\prec)$ contains an antichain of size at least $\sqrt{k}$. First, suppose $(S,\prec)$ contains a chain $v_{1}\prec\cdots\prec v_{\ell}$ of length $\ell\geq\sqrt{k}$. Then $(v_{1},\ldots,v_{\ell})$ is a free set since it is a subsequence of $S$. For each $i\in\{1,\ldots,\ell\}$, let $(x_{i},y_{i})$ be the coordinates given to $v_{i}$ in $\Gamma$. By the definition of free set, $G$ has a straight-line crossing-free drawing in which $v_{i}$ is mapped to $(x_{i},y_{i})$, for each $i\in\{1,\ldots,\ell\}$. Therefore, $\operatorname{fix}(\Gamma)\geq\ell\geq\sqrt{k}$.
To deal with the case where $(S,\prec)$ contains an antichain of size at least $\sqrt{k}$, we use the fact that the ordered set obtained by reversing the order of $S$ is also a free set. Let $(S,\prec^{\prime})$ be the partial order on the elements of $S$ in which $v\prec^{\prime}w$ if and only if $v$ appears after $w$ in $S$ and the ${x}$-coordinate of $v$ in $\Gamma$ is less than the ${x}$-coordinate of $w$ in $\Gamma$. The elements of any antichain in $(S,\prec)$ form a chain in $(S,\prec^{\prime})$. Now we proceed as in the previous paragraph to show that $\operatorname{fix}(\Gamma)\geq\sqrt{k}$. Therefore $\operatorname{fix}(\Gamma)\geq\sqrt{k}$ for any straight-line drawing $\Gamma$ of $G$. Therefore $\operatorname{fix}(G)\geq\sqrt{k}$. ∎
For most classes $\mathcal{X}$ of planar graphs, including trees, outerplanar graphs, and all planar graphs, the best known lower bounds (and for some classes asymptotically optimal bounds) for $\operatorname{fix}_{\mathcal{X}}(n)$ can be obtained by an application of Lemma 1, along with a result on free sets in graphs in the class $\mathcal{X}$ [39, 24, 17, 50]. The sole exception is the class $\mathcal{C}$ of cycles, for which Cibulka [23] uses repeated applications of the Erdős-Szekeres Theorem to obtain a lower bound of $\Omega(n^{2/3})$.
@@
    $S$ is an (ordered) proper-good set.
@@
    $S$ is an (ordered) free-collinear set.
@@
    $S$ is an (ordered) free set.
Note that, unlike the other three definitions, the definition of proper-good set is purely combinatorial/topological; it has no mention of straight-lines or vertex coordinates. In this respect, 1 plays a role similar to Fáry’s Theorem. Using the proper-good set definition allows purely combinatorial methods to be used in the search for free sets. For any proper-good set, there is a short easily-verifiable certificate (a crossing-free drawing of $G$ and a description of the curve $C$) that $S$ is indeed a proper-good set. In contrast, in order to show directly that $S$ is a free-collinear set (or a free set) involves proving a fact that must hold for all choices of $x_{1},\ldots,x_{k}$ (or $(x_{1},y_{1}),\ldots,(x_{k},y_{k})$, respectively). Working with this definition directly would involve (at the very least) a significant amount of linear algebra.
The proof of 1 is spread across a few different works [24, 28, 17], which we describe below. First, we remark that it is immediate from the definitions that every free set is a free-collinear set and every free-collinear set is a collinear set. It is also easy to see that every collinear set is a proper-good set by taking a straight-line crossing-free drawing $\Gamma$ of $G$ with the vertices of the (ordered) collinear set $S$ on the ${x}$-axis in increasing order of ${x}$-coordinate. Let $s$ be a line segment that is contained in the ${x}$-axis and that contains every intersection between $\Gamma$ and the ${x}$-axis in its interior. Then $s$ is proper with respect to $\Gamma$. Finally, make $s$ into a closed curve by joining the endpoints of $s$ with a curve $C$ that completely avoids $\Gamma$.
@@
The proof that every proper-good set is a collinear set is due to Da Lozzo et al. [24] and is illustrated in Fig. 1. For each edge of $G$ that crosses $C$, the proof introduces a dummy vertex (white vertices in Fig. 1(a)) that splits the edge into two edges with a common vertex on $C$. The resulting graph $G^{\prime}$ then has three types of vertices: a set $X$ of inner vertices (contained in the interior of $C$), a set $Y$ of boundary vertices (contained in $C$) and a set $Z$ of outer vertices (the vertices in the exterior of $C$). These steps are depicted in Fig. 1(a).
Refer to caption Refer to caption Refer to caption